# Reinforced In-Context Black-Box Optimization

### v1

@@ 5 Conclusion
Potential Broader Impact
This paper presents work whose goal is to advance the field of black-box optimization. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Model Details
We employ the commonly used GPT architecture [47] and the hyper-parameters are maintained consistently across the problems. Details can be found in Table 1. For BC, BC Filter and OptFormer, the hyper-parameters are set as same as that of our model.
@@ B.1 Benchmarks
  • BBOB [15] is a widely used synthetic BBO benchmark, consisting of $24$ synthetic functions in continue domain. This benchmark makes a series of transformations in the search space, such as linear transformations (e.g., translation, rotation, scaling) and non-linear transformations (e.g., Tosz, Tasy), to create a distribution of functions while retaining similar properties. According to the properties of functions, these synthetic functions can be divided into $5$ categories, i.e., (1) separable functions, (2) moderately conditioned functions, (3) ill-conditioned and unimodal functions, (4) multi-modal functions with adequate global structure, and (5) multi-modal functions with weak global structures. We select one function from each category to evaluate our algorithm, Rastrigin from (1), Rosenbrock Rotated from (2), Sharp Ridge from (3), Greiwank Rosenbrock from (4), and Lunacek from (5). We use the BBOB benchmark implementation in Open Source Vizier. The dimension is set to $10$ for all functions.
  • SVM and XGBoost. HPO-B [2] is the most commonly used HPO benchmark and grouped by search space id. Each search space id corresponds to a machine learning model, e.g., SVM or XGBoost. Each such search space has multiple associated dataset id, which is a particular HPO problem, i.e., optimizing the performance of the corresponding model on a dataset. For continue domain, it fits an XGBoost model as the objective function for each HPO problem. These datasets for each search space id are divided into training and test datasets. We examine our method on two selected search space id, i.e., $5527$ and $6767$, which are two representative HPO problems tuning SVM and XGBoost, respectively. SVM has $3$ parameters, while XGBoost has $18$ parameters, which is the most parameters in HPO-B.
  • Rover Trajectory Planning [57; 16] is a trajectory optimization task to emulate a rover navigation task. The trajectory is determined by fitting a B-spline to $30$ points in a $2$D plane, totally $60$ parameters to optimize. Given $\bm{x}\in[0,1]^{60}$, the objective function is $f(\bm{x})=c(\bm{x})+\lambda(\|\bm{x}_{0,1}-\bm{x}_{s}\|_{1}+\|\bm{x}_{58,59}-\bm{x}_{g}\|_{1})+b$, where $c(\bm{x})$ is the cost of a given trajectory, $\bm{x}_{s}$ and $\bm{x}_{g}$ are $2$D points that specify the starting and goal positions in the plane, $\lambda$ and $b$ are parameters to define the problem. This problem is non-smooth, discontinuous, and concave over the first two and last two dimensions. To construct the distribution of functions, we applied translations in $[-0.1,0.1]^{d}$ as well as scalings in $[0.9,1.1]$, similar to [56; 15]. The training and test tasks are randomly sampled from the distribution.
@@ B.2 Bahavior Algorithms
The datasets are generated with some representatives of heuristic search, EA, and BO as behavior algorithms. We use the implementation in Open Source Vizier [51] for Random Search, Shuffled Grid Search, Eagle Strategy, CMA-ES. For Hill Climbing and Regularized Evolution, we provide a simple reimplementation. We use the implementation in BoTorch [8] for GP-EI. The details of behavior algorithms are summarized as follows.
  • Random Search uniformly at random selects a point in the domain at each iteration.
  • Shuffled Grid Search discretizes the ranges of real parameters into $100$ equidistant points and then selects without replacement a random point from the grid at each iteration.
  • Hill Climbing. At each iteration $t$, the current best solution $x_{best}$ is mutated (using the same mutation operation as Regularized Evolution) to $x_{next}$ and evaluated. If $f(x_{next})\textgreater f(x_{best})$, we update $x_{best}$ to $x_{next}$.
  • Regularized Evolution [49] is an evolutionary algorithm with tournament selection and age-based replacement. We use a population size of $25$ and tournament size of $5$. At each iteration, we randomly select a tournament subset from the current population, and mutate the solution with maximum value. The mutation operation is to uniformly at random selects one of the parameters and mutates to a random value in the domain.
  • Eagle Strategy [63] without the Levy random walk, aka Firefly Algorithm, maintains a population of fireflies. Each firefly emits a light whose intensity corresponds to the objective value. At each iteration, for a firefly, a weight is calculated to chase after a brighter firefly and actively move away from darker one in its vicinity, and the position is updated using the calculated weight.
  • CMA-ES [26] is a popular evolutionary algorithm. At each iteration, the candidate solutions are sampled from a multivariate normal distribution and evaluated, then the mean and covariance matrix are updated.
@@ B.3 Data
For training, a set of tasks are sampled from the task distribution and above behavior algorithms are used to collect data. Specifically, for BBOB suite, a total of $200$ functions are sampled, and each behavior algorithm is executed $500$ times except GP-EI, which is limited to $100$ function samples due to high time cost. For HPO-B, we use the “v3” meta-training/test splits, which consist of $51$ training and $6$ test tasks for SVM, and similarly, $52$ training and $6$ test tasks for XGBoost. All behavior algorithms employ $500$ random seeds in the execution except GP-EI, which is executed using $500$ random seeds. For rover problem, $300$ functions are sampled with each being run $500$ times, whereas for GP-EI, a smaller number of $50$ functions is selected, each being run $500$ times. For testing, we randomly sample $10$ functions for each problem and run each algorithm multiple times to report the average performance and standard deviation.
@@ Appendix C Re-Implementation of OptFormer
OptFormer [14] is a general optimization framework based on transformer [54] and provides an interface to learn policy or function prediction. When provided with textual representations of historical data, OptFormer is capable of determining the next query point $\bm{x}_{t}$ as policy. Additionally, if the histories incorporate a possible query point $\bm{x}_{t}$, the framework is designed to predict the corresponding $y_{t}$, thereby serving as function prediction. We concentrate on the aspect of policy learning, attributable to its greater relevance to our work.
The original implementation is built upon text-based transformer and uses private datasets for training. In this paper, we have re-implemented a simplified version of OptFormer, where we omit the textual tokenization process and only maintain the algorithm type within the metadata. Numerical inputs are fed into the model and we employ different initial states $\bm{x}_{0}$ and $y_{0}$ to represent different algorithms. The hyper-parameters are set as same as our method.
To examine the algorithm imitation capability, we compare our re-implemented with the corresponding behavior algorithms. The results are shown in Figure 4. For the sake of clarity in the visual representation, we only plot a subset of behavior algorithms, including Shuffled Grid Search, Hill Climbing, Regularized Evolution and Eagle Strategy. These behavior algorithms are plotted by solid lines, while their OptFormer counterparts are shown by dashed lines with the same color. Note that the figure is shown using the immediate function values as the $y$-axis to facilitate a comprehensive observation of the optimization.
@@
Appendix D Cross-Distribution Generalization
We conduct experiments to examine the generalization of our method to unseen function distributions. We train the model on $4$ of $5$ chosen synthetic functions and test it with the remaining one. The results in Figure 5 have shown that RIBBO has a strong generalization ability to unseen function distributions.
Figure 5: Cross-Distribution Generalization by training on $4$ synthetic functions and testing on the remaining one.
@@ Appendix E Ablation Studies
We provide further ablation studies to examine the influence of the important components and hyper-parameters of RIBBO, including the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for the function value $y$, the model size, and the sampled subsequence length $\tau$ during training.
@@
Token Aggregation aims to aggregate the information of $(\bm{x}_{i},y_{i},R_{i})$ tokens and establish associations between them. RIBBO employs the Concatenation method (Concat), i.e., concatenating to aggregate $\bm{x}_{i}$, $y_{i}$ and $R_{i}$ to form a single token. It is compared with two different token aggregation methods, i.e., Addition (Add) and Interleaving (Interleave). The addition method integrates the values of each token to form a single token, while interleaving method addresses each token sequentially. The results are shown in Figure 6(a). Concatenation method surpasses both the addition and interleaving methods, with the addition method exhibiting the least efficacy. The concatenation method employs a relatively straightforward technique to aggregate each token by direct concatenation, while the interleaving method adopts a more complex process to learn the interrelations of tokens. The inferior performance of the addition method is because that the summation will omit some details of the original tokens, which are necessary for the generation of the next query $\bm{x}_{t}$.
@@
Normalization Method is to balance the scales of function value $y$ across tasks. We compare the employed random normalization with the dataset normalization and no normalization. Dataset normalization scales the value $y$ by $(y-y_{\rm min}^{i})/(y_{\rm max}^{i}-y_{\rm min}^{i})$ for task $i$, where $y_{\rm min}^{i}$ and $y_{\rm max}^{i}$ denote the observed minimum and maximum values for $f_{i}$, while no normalization entails utilizing the original function values. The results are shown in Figure 6(b). Random normalization and dataset normalization have similar performance, while no normalization leads to the poor performance. The normalization method has a direct effect on the calculation of RTG tokens and thereby influences the generating with a desired regret. No normalization results in significant variances in the scale of $y$, posing a challenge for the generation process. Both random and dataset normalization scale the value of $y$ within a reasonable range, thus facilitating the training and inference. Nevertheless, random normalization can make a model exhibit invariance across various scales of $y$ as mentioned in [61; 14]. Consequently, we recommend to use random normalization in practice.
@@
Model Size has an effect on the in-context learning ability. We compare the performance of the employed model size with a smaller one (8 layers, 4 attention heads and 128-dimensional embedding space) and a larger one (16 layers, 12 attention heads and 384-dimensional embedding space). The results are shown in Figure 6(c), indicating that the model’s size has small impact on its overall performance. A smaller model, with limited capacity, exhibits a diminished capability to the in-context learning ability, while a larger model requires more training data, potentially leading to a slight decrement in performance.
@@
Subsequence Length $\tau$ controls the length of context during both training and inference phases. Employing subsequence sampling can be viewed as a form of data augmentation that facilitates the training process. As shown in Figure 6(d), sampling subsequences instead of using the entire history as context, specifically when $\tau=T=150$, leads to an improvement of the performance. Causal transformer training and inference are quadratic in the context length, hence, a smaller $\tau$ can result in reduced computational complexity. However, an excessively shortened context may contain insufficient historical information, which could, in turn, adversely affect the performance.

### v2

@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2024/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
@@ Appendix A Model Details
We employ the commonly used GPT architecture [48] and the hyper-parameters are maintained consistently across the problems. Details can be found in Table 1. For BC, BC Filter and OptFormer, the hyper-parameters are set as same as those of our model. The training takes about $20$ hours on $1$ GPU (Nvidia RTX $4090$).
@@ B.1 Benchmarks
  • BBOB [15] is a widely used synthetic BBO benchmark, consisting of $24$ synthetic functions in the continuous domain. This benchmark makes a series of transformations in the search space, such as linear transformations (e.g., translation, rotation, scaling) and non-linear transformations (e.g., Tosz, Tasy), to create a distribution of functions while retaining similar properties. According to the properties of functions, these synthetic functions can be divided into $5$ categories, i.e., (1) separable functions, (2) moderately conditioned functions, (3) ill-conditioned and unimodal functions, (4) multi-modal functions with adequate global structure, and (5) multi-modal functions with weak global structures. We select one function from each category to evaluate our algorithm, Rastrigin from (1), Rosenbrock Rotated from (2), Sharp Ridge from (3), Greiwank Rosenbrock from (4), and Lunacek from (5). We use the BBOB benchmark implementation in Open Source Vizier [^1] https://github.com/google/vizier, and the dimension is set to $10$ for all functions.
  • SVM and XGBoost. HPO-B [2] is the most commonly used HPO benchmark and is grouped by search space id. Each search space id corresponds to a machine learning model, e.g., SVM or XGBoost. Each such search space has multiple associated dataset id, which is a particular HPO problem, i.e., optimizing the performance of the corresponding model on a dataset. For the continuous domain, it fits an XGBoost model as the objective function for each HPO problem. These datasets for each search space id are divided into training and test datasets. We examine our method on two selected search space id, i.e., $5527$ and $6767$, which are two representative HPO problems tuning SVM and XGBoost, respectively. SVM has $3$ parameters, while XGBoost has $18$ parameters, which is the most in HPO-B. We use the official implementations [^2] https://github.com/releaunifreiburg/HPO-B.
  • Rover Trajectory Planning [16, 60] is a trajectory optimization task designed to emulate a rover navigation task. The trajectory is determined by fitting a B-spline to $30$ points in a $2$D plane, resulting in a total of $60$ parameters to optimize. Given $\bm{x}\in[0,1]^{60}$, the objective function is $f(\bm{x})=c(\bm{x})+\lambda(\|\bm{x}_{0,1}-\bm{s}\|_{1}+\|\bm{x}_{58,59}-\bm{g}\|_{1})+b$, where $c(\bm{x})$ is the cost of the given trajectory, $\bm{x}_{0,1}$ denotes the first and second dimensions of $\bm{x}$ (similarly for $\bm{x}_{58,59}$), $\bm{s}$ and $\bm{g}$ are $2$D points that specify the starting and goal positions in the plane, $\lambda$ and $b$ are parameters to define the problem. This problem is non-smooth, discontinuous, and concave over the first two and last two dimensions. To construct the distribution of functions, we applied translations in $[-0.1,0.1]^{d}$ and scalings in $[0.9,1.1]$, similar to previous works [15, 59]. The training and test tasks are randomly sampled from the distribution. We use the standard implementation for the rover problem [^3] https://github.com/uber-research/TuRBO.
@@ B.2 Bahavior Algorithms
The datasets are generated with several representatives of heuristic search, EA, and BO as behavior algorithms. We use the implementation in Open Source Vizier [52] for Random Search, Shuffled Grid Search, Eagle Strategy, CMA-ES. For Hill Climbing and Regularized Evolution, we provide a simple re-implementation. We use the implementation in BoTorch [8] for GP-EI. The details of these behavior algorithms are summarized as follows.
  • Random Search selects a point uniformly at random from the domain at each iteration.
  • Shuffled Grid Search discretizes the ranges of real parameters into $100$ equidistant points and selects a random point from the grid without replacement at each iteration.
  • Hill Climbing. At each iteration $t$, the current best solution $x_{best}$ is mutated (using the same mutation operation as Regularized Evolution) to generate $x_{next}$, which is then evaluated. If $f(x_{next})\textgreater f(x_{best})$, $x_{best}$ is updated to $x_{next}$.
  • Regularized Evolution [50] is an evolutionary algorithm with tournament selection and age-based replacement. We use a population size of $25$ and a tournament size of $5$. At each iteration, a tournament subset is randomly selected from the current population, and the solution with the maximum value is mutated. The mutation operation uniformly selects one of the parameters and mutates it to a random value within the domain.
  • Eagle Strategy [66] without the Levy random walk, aka Firefly Algorithm, maintains a population of fireflies. Each firefly emits light whose intensity corresponds to the objective value. At each iteration, for each firefly, a weight is calculated to chase after a brighter firefly and actively move away from darker ones in its vicinity. The position is updated based on the calculated weight.
  • CMA-ES [26] is a popular evolutionary algorithm. At each iteration, candidate solutions are sampled from a multivariate normal distribution and evaluated, then the mean and covariance matrix are updated.
@@ B.3 Data
For training, a set of tasks is sampled from the task distribution and the above behavior algorithms are used to collect data. Specifically, for the BBOB suite, a total of $200$ functions are sampled, and each behavior algorithm is executed $500$ times, except for GP-EI, which is limited to $100$ function samples due to its high time cost. For HPO-B, we use the “v3” meta-training/test splits, which consist of $51$ training and $6$ test tasks for SVM, and similarly, $52$ training and $6$ test tasks for XGBoost. All behavior algorithms employ $500$ random seeds in the execution. For the rover problem, $300$ functions are sampled with each being run $500$ times, whereas for GP-EI, a smaller number of $50$ functions is selected, each being run $500$ times. For testing, we randomly sample $10$ functions for each problem and run each algorithm multiple times to report the average performance and standard deviation.
@@ Appendix C Re-Implementation of OptFormer
OptFormer [14] is a general optimization framework based on transformer [57] and provides an interface to learn policy or function prediction. When provided with textual representations of historical data, OptFormer can determine the next query point $\bm{x}_{t}$, acting as a policy. Additionally, if the context incorporate a possible query point $\bm{x}_{t}$, the framework is capable of predicting the corresponding $y_{t}$, thus serving as a prediction model. We focus on the aspect of policy learning, due to its greater relevance to our work.
The original implementation is built upon a text-based transformer and uses private datasets for training. In this paper, we have re-implemented a simplified version of OptFormer, where we omit the textual tokenization process and only retain the algorithm type within the metadata. Numerical inputs are fed into the model and we use different initial states $\bm{x}_{0}$ and $y_{0}$ to distinguish among algorithms. The initial states are derived by indexing the algorithm type in an embedding layer, thereby enabling the identification of distinct behavior algorithms within OptFormer. The hyper-parameters are set as same as our method.
To examine the algorithm imitation capability, we compare our re-implementation with the corresponding behavior algorithms. The results are shown in Figure 4. For clarity in the visual representation, we only plot a subset of the behavior algorithms, including Shuffled Grid Search, Hill Climbing, Regularized Evolution and Eagle Strategy. These behavior algorithms are plotted by solid lines, while their OptFormer counterparts are shown by dashed lines with the same color. Note that the figure uses the immediate function values as the $y$-axis to facilitate a comprehensive observation of the optimization process.
@@
Appendix D Generalization
To examine the generalization of our method to unseen function distributions, we conduct experiments to train the model on $4$ of $5$ chosen synthetic functions and test on the remaining one. The results in Figure 5 have shown that RIBBO has a strong generalization ability to unseen function distributions.
Figure 5: Cross-distribution generalization by training on $4$ of $5$ chosen synthetic functions and testing on the remaining one.
We also train the model across all training datasets from BBOB synthetic functions simultaneously, and evaluate on the same functions used in previous sections, specifically Greiwank Rosenbrock, Lunacek, Rastrigin, Rosenbrock, and Sharp Ridge. These functions serve as representatives of each category within the BBOB suite. The results are shown in Figure 6. RIBBO demonstrates relatively inferior performance compared to some behavior algorithms, which might be attributed to that the training dataset consists of excessive trajectories from diverse behavior algorithms across different functions, and complicates the process of algorithm identification solely through RTG tokens. Song et al. [53] had similar observation in Section 5.2 of their paper, indicating the significance of textual cues for $\bm{x}$ and the complete failure of the model when using anonymized inputs. However, RIBBO still surpasses both BC and BC Filter, emphasizing the effectiveness of RTG tokens. RIBBO employs RTG tokens to facilitate the identification of diverse algorithms across various functions and to generate sequences with specified regret. Further improvement could be achieved by integrating textual representations for $\bm{x}$ and metadata describing the problem, which may potentially improve the model’s ability to identify and generalize [54]. This leaves as future work.
Figure 6: Generalization by training across the entire BBOB suite.
@@ Appendix E Ablation Studies
We provide further ablation studies to examine the influence of the key components and hyper-parameters in RIBBO, including the method to aggregate $(\bm{x}_{i},y_{i},R_{i})$ tokens, the normalization method for the function value $y$, the model size, and the sampled subsequence length $\tau$ during training.
@@
Token Aggregation aims to aggregate the information from $(\bm{x}_{i},y_{i},R_{i})$ tokens and establish associations between them. RIBBO employs the Concatenation method (Concat), i.e., concatenating to aggregate $\bm{x}_{i}$, $y_{i}$ and $R_{i}$ to form a single token. This method is compared with two alternative token aggregation methods, i.e., Addition (Add) and Interleaving (Interleave). The addition method integrates the values of each token into one, while the interleaving method addresses tokens sequentially. The results are shown in Figure 7(a). The concatenation method surpasses both the addition and interleaving methods, with the addition method showing the least efficacy. The concatenation method employs a relatively straightforward approach to aggregate tokens, while the interleaving method adopts a more complex process to learn the interrelations among tokens. The inferior performance of the addition method can be due to the summation operator, which potentially eliminates critical details from the original tokens. These details are essential for generating the subsequent query $\bm{x}_{t}$.
@@
Normalization Method is to balance the scales of the function value $y$ across tasks. We compare the employed random normalization with the dataset normalization and the absence of any normalization. Dataset normalization adjusts the value of $y$ by $(y-y_{\rm min}^{i})/(y_{\rm max}^{i}-y_{\rm min}^{i})$ for each task $i$, where $y_{\rm min}^{i}$ and $y_{\rm max}^{i}$ denote the observed minimum and maximum values of $f_{i}$. No normalization uses the original function values directly. The results in Figure 7(b) indicate that random normalization and dataset normalization perform similarly, whereas the absence of normalization significantly hurts the performance. The choice of normalization method directly impacts the computation of RTG tokens and subsequently affects the generation of the desired regret. The lack of normalization leads to significant variations in the scale of $y$, which complicates the training process. Both random and dataset normalization scale the value of $y$ within a reasonable range, thus facilitating both training and inference. However, random normalization can bring additional benefits, such as invariance across various scales of $y$ as mentioned in Wistuba and Grabocka [64] and Chen et al. [14]. Therefore, we recommend using random normalization in practice.
@@
Model Size has an effect on the in-context learning ability. We assess the effects of different model sizes by comparing the performance of the currently employed model size with both a smaller and a larger model. The smaller model consists of $8$ layers, $4$ attention heads, and $128$-dimensional embedding space, while the larger model has $16$ layers, $12$ attention heads, and $384$-dimensional embedding space. The results are shown in Figure 7(c), indicating that the model size has a minimal impact on overall performance. Specifically, the smaller model, due to its limited capacity, shows a reduced performance, while the larger one, potentially more powerful, requires more training data, which can lead to a slight decrease in performance due to overfitting or inefficiency in learning from a limited dataset. This analysis highlights the trade-offs involved in selecting the appropriate model size for optimal performance.
@@
Subsequence Length $\tau$ controls the context length during both the training and inference phases. The process of subsequence sampling acts as a form of data augmentation, enhancing training efficiency. As shown in Figure 7(d), sampling subsequences rather than using the entire history as context, particularly when $\tau=T=150$, leads to improved performance. The computational complexity for causal transformer training and inference scales quadratically with the context length. Therefore, utilizing a shorter $\tau$ can significantly reduce computational demands. However, it is important to note that a shorter $\tau$ might not capture sufficient historical data, potentially degrading the performance due to the insufficient contextual information. This highlights the trade-offs between computational complexity and performance.